If 5/6 is divided by a certain fraction, the result is 2/5. What is the fraction?
Answer by ReadingBoosters(3246)

5/6 / x = 2/5 which is the same as 5/6*1/x=2/5
5/6x=2/5
5=6x(2/5)
5=12/5x
5 * 5/12 = x
x = 25/12 or 2 1/12
proof
5/6 / 25/12 = 2/5
..........................
Delighted to help
-Reading Boosters